A shipping crate is to be constructed with its length six times the size of its width, and its height one-third the size of its length. The volume of the box must be 6144 cubic inches. Find the dimensions of the box.

thanks.

Answer by checkley77(12844)

L=6W
H=L/3
LWH=6144
6W*W*L/3=6144
6W*W*6W/3=6144
36W^3/3=6144
12W^3=6144
W^3=6144/12
W^3=512
W=CUBERT512
W=8 FOR THE WIDTH
L=6*8=48 FOR THE LENGTH
H=48/3=16 FOR THE HEIGHT.
PROOF:
8*48*16=6144
6144=6144